6x^2-x-2=0 help!!!!
solve the following equations

Answer by checkley79(3341)

6x^2-x-2=0
(3x-2)(2x+1)=0
3x-2=0
3x=2
x=2/3 ans.
2x+1=0
2x=-1
x=-1/2 ans.